Two cars A and B are moving along a straight road in the same direction with velocities of 25 km/h and 40 km/h, respectively. Find the velocity of car B relative to car A.


1
Expert's answer
2021-03-29T09:03:47-0400

Relative velocity of B with respect to A = V(B) - V(A) = 40 - 25 = 15 km/h